Suppose that the mean stopping time of a particular model of car moving at 40 mph is listed

as 1.8 seconds, with a standard deviation of 0.25 seconds. A random sample of 12 cars obtains a

standard deviation of 0.35 seconds. Test the claim that o = 0.25 seconds at the 0.05 level of significance


1
Expert's answer
2022-05-25T17:25:44-0400

The following null and alternative hypotheses need to be tested:

"H_0:\sigma^2=0.25^2=0.0625"

"H_a:\sigma^2\not=0.25^2=0.0625"

This corresponds to a two-tailed test test, for which a Chi-Square test for one population variance will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=11" degrees of freedom, and the the rejection region for this two-tailed test is "R = \{\chi^2: \chi^2 < 3.8157 \text{ or } \chi^2 > 21.92\}."

The Chi-Squared statistic is computed as follows:


"\chi^2 =\dfrac{(n-)s^2}{\sigma^2}=\dfrac{(12-1)(0.35)^2}{(0.25)^2}=21.56"

Since it is observed that "\chi_L^2 = 3.8157 \le \chi^2 = 21.56 \le \chi_U^2 = 21.92,"

it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma^2" is different than "(0.25)^2," at the "0.05" significance level.